OpenStudy (skylyerbailey):

If the slope is 1/3 and a point on the line is (6,1), what is the equation of the line?

OpenStudy (jiteshmeghwal9):

y-y_1=m(x-x_1) (x_1,y_1)=(6,1) m=1/3

OpenStudy (jiteshmeghwal9):

find the equation now :)